OpenStudy (anonymous):

Which of the following statements uses the commutative property of addition? 2 + (3 + 4) = (2 + 3) + 4 2 + (3 + 4) = (2 + 3) + (2 + 4) 2 + (3 + 4) = 2 + (4 + 3)

OpenStudy (anonymous):

Do you know the definition of commutative?

OpenStudy (anonymous):

It means that when performing an operator of some sort, the order in which you do it in does not matter.

OpenStudy (anonymous):

But if the order does not matter then all of them are a possible answer?

OpenStudy (anonymous):

Ahh here is where you have to be careful :P . The Brackets makes a difference.

OpenStudy (anonymous):

Oh I see. So it would be the last answer, right?

OpenStudy (anonymous):

2 + (3 + 4) = (2 + 3) + 4 We got 3+4 and then add 2. Or we got 2+3 and add 4. We are essentially taking the same sum but in a different order. 2 + (3 + 4) = (2 + 3) + (2 + 4) We got 3+4 and then add 2. However, we do not perform the same order that we performed on the first case. 2 + (3 + 4) = 2 + (4 + 3) Same as the first.

OpenStudy (anonymous):

1st and last actually because the first 1 uses the same idea.

OpenStudy (anonymous):

But I can only choose one answer. So The first one would be the one I should lean more towards?

OpenStudy (anonymous):

Last one is more "better" actually.

OpenStudy (anonymous):

Wait!!!

OpenStudy (anonymous):

Nvm, the first one... The order changes.

OpenStudy (anonymous):

I am a bit slow today.

OpenStudy (anonymous):

Because you perform the same thing in a different order but the both produce the same result.

OpenStudy (anonymous):

Thats what I was thinking when you said the parenthesis matter. Thank you so much @Dido525 .

OpenStudy (anonymous):

For the 3rd one, 3+4 and 43 is the same thing but it's not performed in a different order.

OpenStudy (anonymous):

4+3*

OpenStudy (anonymous):

yeah not a problem :) .